OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively.

OpenStudy (anonymous):

Geometric sequences are like: a , a*r, a*r*r, a*r*r*r so on.. \[G _{5} = G _{2}*r ^{3}\] Then \[r ^{3} = G _{5}/G _{2} = 768/-12 = -64 \rightarrow r = -4\] \[-12 = G _{2} = r*G _{1} = r*a = -4*a \rightarrow a=3\] So our constant is 3 then nth term would be\[G _{n} = a*r ^{n} = 3*(-4)^{n}\]

OpenStudy (anonymous):

Here are my choices: an = 3 • (-4)n + 1 an = 3 • 4n - 1 an = 3 • (-4)n - 1 an = 3 • 4n

OpenStudy (anonymous):

Those answers are wrong :/

OpenStudy (anonymous):

Those are my options...

OpenStudy (anonymous):

Sorry, its my bad. the general term for geometric series is:\[G _{n} = a*r ^{n-1}\] Because the first term is not a*r but a. So power of r should be 0 for n=1. \[G _{n} = 3*(-4)^{n-1}\]

OpenStudy (anonymous):

Thank You !